# Equivariant Spherical Transformer for Efficient Molecular Modeling

### v2

@@ 5 Conclusion
We present EST, a new SE(3)-equivariant framework for modeling molecules. By integrating a Transformer structure with steerable vectors, EST offers greater expressive power than tensor-product-based frameworks. We showed EST’s strong performance on the OC20 and QM9 datasets. A current limitation is the equivariance loss caused by spherical sampling. Future work will focus on improving spherical sampling uniformity. This could lead to fewer sampling points and even better equivariance.
@@ A.1.2 Fourier transformation over $\mathbb{S}^{2}$
Using the fact $\mathbf{Y}^{l}(\mathbf{R}\vec{\mathbf{p}})=\mathbf{D}^{l}(\mathbf{R})\mathbf{Y}^{l}(\vec{\mathbf{p}}),$ and Equation (20), we know
@@ A.2.1 Equivariance of Clebsch-Gordan Tensor Product
Here, we omit the input argument of the spherical harmonics, which can represent any direction on the sphere. Equation (26) represents a relationship between scalar. If we transform the scalar to vector or matrix like what we do in Equation (3), Equation (26) is equal to
@@ B.1 Proof for Expressiveness of EST
Recall that CG coefficients are in fact the expansion coefficients of a product of two spherical harmonics in terms of a single spherical harmonic (see Equation (17)):
$Y^{(l_{1},m_{1})}Y^{(l_{2},m_{2})}=$ $\sum_{l_{0},m_{0}}\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)}{4\pi(2l_{0}+1)}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(0,0\right)}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}Y^{l_{0},m_{0}}$ (30)
Equation (LABEL:equ:mul_u_v) can be transformed to
where $\mathbf{H}^{(l_{0},m_{0})}_{(l_{1},m_{1}),(l_{2},m_{2})}=\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)}{4\pi(2l_{0}+1)}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(0,0\right)}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}$. Compared to $\mathbf{C}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$, the term $\mathbf{H}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$ introduces an additional constant term that can be approximated by linear layers and the FFNs. Therefore, Equation (LABEL:equ:mul_u_v) is equivalent to Equation (28). On the other hand, Equation (LABEL:equ:mul_u_v) is inherently part of the Transformer architecture: when relationships between different orientations are masked and the $\mathbf{V}$ vectors are taken as constant vectors, the Transformer can naturally reduce to Equation (LABEL:equ:mul_u_v). ∎
Through the above proof, we find that the spherical representation of the output from the CG tensor product is fundamentally a special case of EST. Furthermore, EST can capture dependencies between different orientations, which is particularly beneficial for approximating higher degree spherical harmonic representations. For instance, at a direction $(\theta,\varphi)$, higher-degree spherical harmonics may involve contributions from other directions such as $(2\theta,\varphi),(\theta,2\varphi),(2\theta,2\varphi)$ (see $Y^{(2)}$ in Equation (19)). This property also enables EST to approximate equivariant features of higher degree than the input.
B.2 Proof for Equivariance of EST
The proof for Theorem 2 is shown in the following.
Proof.
@@ B.2 Proof for Equivariance of EST
where $a(\cdot)$ denotes the attention coefficients. Terms $Q,K,V$ denote three independent linear transformations. When the origin representation $\mathbf{x}$ is rotation by Wigner-D matrix $\mathbf{D}^{-1}$, the representation $f(\vec{\mathbf{p}})$ is transformed to $f(\mathbf{R}\vec{\mathbf{p}})$, where $\mathbf{R}$ and $\mathbf{D}$ share the same transformation parameters (see Equation (23)). Therefore, the attention results are changed to
@@
In the implementation, the continuous function must be represented by discrete sampling. The attention operation become
@@
After introducing random rotations, Equation (35) becomes
@@
Let us consider an ideal case where the point density is the same across all local regions. That is, the points on the sphere are perfectly uniformly distributed. In this scenario, the points in any region are symmetric, and we can find a set of $\vec{\mathbf{p}}_{3}$ that has a one-to-one correspondence with all $\mathbf{R}\vec{\mathbf{p}}_{2}$, i.e., the difference between the set of points $\vec{\mathbf{p}}_{2}$ and the set of points $\vec{\mathbf{p}}_{3}$ is only in their order. Therefore, spherical attention can still strictly preserve equivariance. In contrast, if we cannot find points that correspond one-to-one to all $\mathbf{R}\vec{\mathbf{p}}_{2}$, equivariance will be compromised. To reduce this equivariance loss, we can employ approximately uniform sampling methods, such as Fibonacci Lattices.
@@ B.3 Equivariance of Relative Orientation Embedding
The equivariance of relative orientation embedding can be proven with the same way in Equation (LABEL:equ:proof_equivariance), where $a\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1}),f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})\big{)}$ is transformed to
@@ C.3 Mixture of Hybrid Experts in EST
C.4 Ablation Studies
@@ D.1 Implementation Details of Baselines
Energy coefficient $\lambda_{E}$ 44 11
@@ D.2 Implementation Details of EST Experiments
Hyper-parameters EST EST (with GA)
@@ D.3 Comparison between GotenNet and EST
Comparing with GotenNet, a variant of EST was designed incorporating the GATA message block and substituting the original updating block with our proposed mixture of hybrid experts. The steerable part of the hybrid expert maintain the architecture of EQFF in GotenNet, while the EST updating module serves as the spherical part used in parallel. To maintain a lightweight design, we simply integrate one steerable expert with one spherical expert. Our model adopts the same number of layers as GotenNet but achieves better performance across 6 tasks (table 11).
To ensure a fair comparison, we use training configurations similar to those of the original GotenNet (see Table 12), including a batch size of $32$ and a training schedule of $1000$ epochs. It is notable that several training hyper-parameters of EST differ from those used of GotenNet. For example, GotenNet specifies 10,000 warm-up steps, whereas our model sets this to 0. Additionally, we employed MAE as the loss function, while MSE was used for GotenNet instead. Properties where EST performs poorly may stem from such subtle differences. For example, EST outperforms GotenNet on both $\varepsilon_{HOMO}$ and $\varepsilon_{LUMO}$, but it underperforms significantly on the $\Delta\varepsilon$ property, which is the gap between $\varepsilon_{HOMO}$ and $\varepsilon_{LUMO}$. We suspect that this might be due to the absence of a warmup stage, causing the model to converge to a suboptimal local minimum.
@@
EST (4 layers with GATA) .030 31.2 16.7 13.3 .0070 .019 5.31 4.02 0.029 3.96 3.94 1.16
Table 12: Hyper-parameters for the EST and GotenNet models on QM9 experiments. Red highlights the configurations that differ.
Hyper-parameters EST (with GATA) GotenNet
@@
Learning rate scheduling reduce on plateau linear warmup with reduce on plateau
@@
Optimizer AdamW AdamW
Maximum learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Batch size $32$ $32$
Max training epochs $1000$ $1000$
Dropout rate $0.1$ $0.1$
Number of RBFs $64$ $64$
Cutoff radius (Å) 55 55
$L_{max}$ 22 22
Number of layers 44 44
Node dimension $256$ $256$
Edge dimension $256$ $256$
Number of attention heads 88 88
Intermediate dimension and the number of $512,1$ $512,1$
steerable FFN
Intermediate dimension and the number of $512,1$ -
spherical FFN
Number of spherical point samples $64$ -

### v3

@@ 5 Conclusion
We present EST, a new SE(3)-equivariant framework for modeling molecules. By integrating a Transformer structure with steerable vectors, EST offers greater expressive power than tensor-product-based frameworks. We showed EST’s strong performance on the OC20 and QM9 datasets. A current limitation is the slight equivariance loss caused by spherical sampling. Future work will focus on improving spherical sampling uniformity. This could lead to better equivariance.
@@ A.1.2 Fourier transformation over $\mathbb{S}^{2}$
Using the fact $\mathbf{Y}^{l}(\mathbf{R}\vec{\mathbf{p}})=\mathbf{D}^{l}(\mathbf{R})\mathbf{Y}^{l}(\vec{\mathbf{p}}),$ and (20), we know
@@ A.2.1 Equivariance of Clebsch-Gordan Tensor Product
Here, we omit the input argument of the spherical harmonics, which can represent any direction on the sphere. (26) represents a relationship between scalar. If we transform the scalar to vector or matrix like what we do in (3), (26) is equal to
@@ B.1 Proof for Equivariance of EST
Recall that FT can be represented by
$f(\vec{\mathbf{p}})=\sum_{l=0}^{L}\sum_{m=-l}^{l}x^{(l,m)}Y^{(l,m)}(\vec{\mathbf{p}})=(\mathbf{x}^{(0\rightarrow L)})^{T}\mathbf{Y}^{(0\rightarrow L)}(\vec{\mathbf{p}}).$ (28)
When we apply a random rotation $\mathbf{D}^{-1}$ to $\mathbf{x}$, we obtain
$(\mathbf{D}^{-1}\mathbf{x}^{(0\rightarrow L)})^{T}\mathbf{Y}^{(0\rightarrow L)}(\vec{\mathbf{p}})$ (29)
$=$ $(\mathbf{x}^{(0\rightarrow L)})^{T}\mathbf{D}\mathbf{Y}^{(0\rightarrow L)}(\vec{\mathbf{p}})$ (30)
$=$ $(\mathbf{x}^{(0\rightarrow L)})^{T}\mathbf{Y}^{(0\rightarrow L)}(\mathbf{R}\vec{\mathbf{p}})$ (31)
$=$ $f(\mathbf{R}\vec{\mathbf{p}}),$ (32)
where $\mathbf{R}$ and $\mathbf{D}$ share the same transformation parameters.
@@
where $a(\cdot)$ denotes the operation to compute attention coefficients. Terms $Q,K,V$ denote three independent linear transformations. When the origin representation $\mathbf{x}$ is rotation by a Wigner-D matrix $\mathbf{D}^{-1}$, the representation $f(\vec{\mathbf{p}})$ is transformed to $f(\mathbf{R}\vec{\mathbf{p}})$. Therefore, the attention results are changed to
@@
Note that (LABEL:equ:proof_equivariance) rely on continuous sphere, i.e closed set of orientations $\vec{\mathbf{p}}$. In the implementation, the continuous function must be represented by discrete sampling. The attention operation become
@@
After introducing random rotations, (36) becomes
@@
Let us consider an ideal case where the point density is the same across all local regions. That is, the points on the sphere are perfectly uniformly distributed. The entire sphere is symmetric. The set of $\vec{\mathbf{p}}_{1}$, $\vec{\mathbf{p}}_{2}$, and $\vec{\mathbf{p}}_{3}$ all come from a set of sampling points. For this set, if one point $\vec{\mathbf{p}}_{i}$ can be mapped to another point $\vec{\mathbf{p}}_{j}$ through a specific rotation $\mathbf{R}$, then the other points can also find their corresponding points through this rotation due to perfect symmetry. In this case, we say that the set of sampling points is closed under a special subset of rotations. The model is strictly equivariant on these special rotations.
@@
Enhanced Uniformity
Building on the FL sampling, we further propose a scattering optimization strategy similar to molecular dynamics. We define a distance-based interaction force between every two points: the greater the distance, the smaller the force. Then, we iterate two processes: 1. Simultaneously optimize the three-dimensional coordinates of all points through the interaction forces; 2. Re-project all points back onto the sphere. After multiple rounds of iteration, the spherical sampling points can become more uniform. It is worth noting that both the FL sampling and the dynamics simulation are conducted during the model initialization phase. Once we obtain the uniform sampling points, we save their results and directly read them in the model’s feedforward process.
B.2 Expressiveness of EST
In Proposition 2, wo claim that the function space spanned by the spherical Transformer encompasses that spanned by the CG tensor product and give an intuitive explanation. Here, we provide a mathematical explanation for further understanding.
Recall that CG coefficients are in fact the expansion coefficients of a product of two spherical harmonics in terms of a single spherical harmonic (see (17)):
$Y^{(l_{1},m_{1})}Y^{(l_{2},m_{2})}=$ $\sum_{l_{0},m_{0}}\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)}{4\pi(2l_{0}+1)}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(0,0\right)}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}Y^{l_{0},m_{0}},$ (40)
(LABEL:equ:mul_u_v) can be transformed to
where $\mathbf{H}^{(l_{0},m_{0})}_{(l_{1},m_{1}),(l_{2},m_{2})}=\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)}{4\pi(2l_{0}+1)}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(0,0\right)}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}$. Compared to $\mathbf{C}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$, the term $\mathbf{H}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$ introduces an additional constant term that can be approximated by linear layers and the FFNs. Therefore, (LABEL:equ:mul_u_v) is equivalent to (38). On the other hand, (LABEL:equ:mul_u_v) is inherently part of the Transformer architecture: when relationships between different orientations are masked and the $\mathbf{V}$ vectors are taken as constant vectors, the Transformer can naturally reduce to (LABEL:equ:mul_u_v).
Through the above process, we find that the spherical representation of the output from the CG tensor product is fundamentally a special case of EST. Furthermore, EST can capture dependencies between different orientations, which is particularly beneficial for approximating higher degree spherical harmonic representations. For instance, at a orientation $(\theta,\varphi)$, higher-degree spherical harmonics may involve contributions from other orientations such as $(2\theta,\varphi),(\theta,2\varphi),(2\theta,2\varphi)$ (see $Y^{(2)}$ in (19)). The Transformer in EST can directly combine these lower-degree terms in these orientations to approximate the higher-order terms at the orientation $(\theta,\varphi)$.
@@ B.3 Equivariance of Relative Orientation Embedding
The equivariance of relative orientation embedding can be easily proven with the same way in (LABEL:equ:proof_equivariance), where $a\big(f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1}),f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})\big)$ is transformed to
@@ C.3 Mixture of Hybrid Experts in EST
C.4 Integration to Equivariant Architectures
EST with MoE can be employed both in the message block to compute edge features and in the update block to refine node embeddings. Figure 3(b,c) defines a message passing layer with EST. When computing the message, the $\mathbf{Q}$ in SA come from the aggregation of node $i$ and $j$ embeddings, while $\mathbf{K}$ and $\mathbf{V}$ is derived from the spherical harmonic representation of the relative position $\vec{\mathbf{r}}_{ij}$. In the SA of the update block, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are all derived from the aggregated message. Additionally, in the message and update blocks shown in Figure 3, the spatial expert is used after spherical attention, and the steerable expert is employed as a parallel branch. They are combined using gate weights.
Figure 3: The architecture and building blocks of EST. SH and DTP denote spherical harmonic embedding and depth-wise tensor product [10], respectively. For simplicity, the Fourier and inverse Fourier transform steps are omitted. (a) Overall architecture. (b) Message block. (c) Update block. (d) Two experts operating on the steerable and spatial representations, respectively.
C.5 Ablation Studies
@@ D.1 Implementation Details of Baselines
Energy coefficient $\lambda_{E}$ $4,2$ 11
@@ D.2 Implementation Details of EST Experiments
Hyper-parameters EST-based model (Figure 3) EST (with GA)
@@ D.3 Comparison between GotenNet and EST
Comparing with GotenNet, a variant of EST was designed incorporating the GATA message block and substituting the original updating block with our proposed mixture of hybrid experts. The steerable part of the hybrid expert maintain the architecture of EQFF in GotenNet, while the EST updating module serves as the spherical part used in parallel. To maintain a lightweight design, we simply integrate one steerable expert with one spherical expert. Our model adopts the same number of layers as GotenNet but achieves better performance across 8 tasks (table 11). In addition, to ensure a fair comparison, we use training configurations similar to those of the original GotenNet (see Table 12), including a batch size of $32$ and a training schedule of $1000$ epochs. It is notable that several training hyper-parameters of EST differ from those used of GotenNet. For example, GotenNet specifies 10,000 warm-up steps, whereas our model sets this to 0. Additionally, we employed MAE as the loss function, while MSE was used for GotenNet instead.
Notablly, these target properties can be broadly categorized into energy-related variables (e.g., $G,H,U,U_{0}$, etc.) and molecular geometry-related variables (e.g., $\alpha,\mu$, etc.). We further observed that these two categories of properties exhibit different convergence characteristics during training, which aligns with previous work [4]. Specifically, the energy-related properties are more sensitive to hyperparameters such as the optimizer and learning rate schedule, while the geometry-related properties are relatively less sensitive. Given this, we divided the properties into two groups and tuned different training hyperparameters for each group (see Table 12). It is worth noting that GotenNet has already achieved low error levels on these properties, meaning the scope for further improvement is quite limited.
@@
EST (4 layers with GATA) .030 31.2 16.7 13.3 .0070 .019 5.40 3.71 0.029 3.70 3.70 1.19
Table 12: Hyper-parameters for the EST (with GATA) models on QM9 experiments.
Hyper-parameters EST (with GATA)
properties $\alpha$, $\Delta\varepsilon$, $\varepsilon_{HOMO}$, $\varepsilon_{LUMO}$, $\mu$, $C_{v}$, $R^{2}$ $G$, $H$, $U$, $U_{0}$, U, ZPVE
@@
Intermediate dimension and the number of steerable FFN 512, 1 512, 4
Intermediate dimension and the number of spherical FFN 512, 1 512, 4
Learning rate scheduling linear warmup with reduce on plateau
Optimizer AdamW
Maximum learning rate $1\times 10^{-4}$
Batch size 32
Max training epochs 1000
Dropout rate 0.1
Number of RBFs 64
Cutoff radius (Å) 5
Lmax 2
Number of layers 4
Node dimension 256
Edge dimension 256
Number of attention heads 8
Number of spherical point samples 64